Sol Summary – November 30th

Crew 183+ Sol Summary Report 30November2017

Sol 3

Title: Further Exploration

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I spent from 10:00 AM to 2:00 PM, exploring the area south of Kissing Camel Range. I’d estimate that I covered anywhere from 6-8 km during that time frame. There were many interesting areas that I found, such as a couple of caves that a person can walk into. Although I didn’t find any dinosaur bones out there, I did find a lot of petrified wood. During this time, Shannon managed to complete the MDRS user manual that she’s been working on.

Between 3:00 PM & 5:00 PM, Shannon and I explored the last of the unchecked roads we still needed to check up on.

Look Ahead Plan:

Tomorrow, Curtis and his team will come and fix the water heater. Shannon and I will clean up the Lab for Crew 184.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: Region South of Kissing Camel Range & Road exploration

Reports to be Filed:

-Operations Report

Support Requested: N/A

Operations Report – November 30th

Crew 183 + Operations Report 30NOV2017

SOL: n/a

Name of person filing report:  Shannon Rupert

Non-nominal systems: Nothing new to report

Notes on non-nominal systems:  n/a

Generator (hours run): off 8 am on 6 pm

Solar— SOC 90% (Before generator is run at night)

Diesel –  50 %

Propane – 100 psi

Ethanol Free Gasoline (5 Gallon containers for ATV) – not known gallons.

Water (trailer) – 0 gallons.

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter:  not known

Toilet tank emptied: no

ATV’s Used: Honda

Oil Added? No

ATV Fuel Used:  2 Gals

# Hours the ATVs were Used today: 4

Notes on ATVs: nothing to report

Deimos rover used:  yes


Beginning charge: 100

Ending charge: 94

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Installation in progress


Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Installation in progress


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:


Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Burn barrel not complete. Turns out you
can’t just poke holes in it with a pickaxe.  We have a plan for
tomorrow.  Also, contractors will be onsite all day tomorrow.

Summary of internet:  not known

Summary of suits and radios: nothing to report

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Door was opened this afternoon.
Thinned the babies.  Used 4 gallons of water for watering and prepping
the soil for herbs.

Summary of ScienceDome operations:  Brandon continues his work.
Nothing more to report.

Summary of RAMM operations: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: Nothing tonight, thank you.

Copyright © The Mars Society. All rights reserved. | Main Site